OpenStudy (anonymous):

what is the average kinetic energy of a gas at 285 kelvin (r = 8.314 j/k-mol)

OpenStudy (anonymous):

as we know that average kinetic of a gas is given by the equation K.E=3(K)(T)/2 where value of k=1.3*10^-23 by putting the value of temperature you can find avg k.e

OpenStudy (fionacndg):

so what is the answer then? 3.55 x 103 J/mol??

OpenStudy (fionacndg):

@jdoe0001 @aaronq

OpenStudy (fionacndg):

answer choices- 2.37 x 103 J/mol 7.11 x 103 J/mol 3.55 x 103 J/mol 1.77 x 103 J/mol

OpenStudy (fionacndg):

@e.mccormick